Please,Solve by substitution.Check the answer as well
4x - 10y=6
4x + y=6

Answer by ewatrrr(10682)


Hi,
Solve by substitution
4x - 10y = 6
4x + y = 6 4x = 6-y
4x - 10y = 6
(6 - y) - 10y = 6
6 - y - 10y = 6
-11y = 0
y = 0
4x + y = 6
4x + 0 = 6
4x = 6
x = 6/4 Or 3/2
checking our answer
4x - 10y = 6
4*(3/2) + 0 = 6
6 + 0 = 6